### v2

# Generation and descent for derived categories of noncommutative structure sheaves

Abstract.
This work demonstrates classical generation is preserved by the derived pushforward along the canonical morphism of a noncommutative structure sheaf to its underlying scheme. There are intriguing examples illustrating this phenomenon, particularly from noncommutative resolutions, categorical resolutions, and homological projective duality. Additionally, we establish that the Krull dimension of a variety over a field is a lower bound for the Rouquier dimension of the bounded derived category associated with a noncommutative structure sheaf on it. This is an extension of a classical result of Rouquier to the noncommutative context.
Key words and phrases:
Derived categories, strong generation, Rouquier dimension, noncommutative structure sheaves, noncommutative crepant resolutions, categorical resolution of singularities, homological projective duality, Azumaya algebra
2020 Mathematics Subject Classification:
14F08 (primary), 14A22, 13D09, 16S38, 16E35, 14A30
1. Introduction
In this note, we examine the behavior of generation along the derived pushforward of the canonical morphism from a noncommutative structure sheaf to its underlying scheme. As a result, our investigation introduces new methods coming from noncommutative homological techniques for establishing upper bounds on the Rouquier dimension within triangulated categories relevant to the commutative world.
Consider a Noetherian scheme $X$. The Rouquier dimension of $D^{b}_{\operatorname{coh}}(X)$, denoted by $\dim D^{b}_{\operatorname{coh}}(X)$, is the smallest integer $n$ such that there exists an object $G$, where the smallest subcategory generated by $G$ using finite direct sums, summands, shifts, and at most $n+1$ cones coincides with $D^{b}_{\operatorname{coh}}(X)$. Any such subcategory is denoted as $\langle G\rangle_{n}$, and objects $G$ satisfying these conditions are called strong generators. More generally, an object $G$ is a classical generator if the smallest triangulated subcategory containing $G$ that is closed under direct summands coincides with $D^{b}_{\operatorname{coh}}(X)$. For further background, see Section 2.
There has been recent advancements that illuminate sufficient conditions under which $D^{b}_{\operatorname{coh}}(X)$ possesses such objects: for strong generators, any Noetherian quasi-excellent separated scheme of finite Krull dimension [Aok21]; for classical generators, any Noetherian $J\textrm{-}2$ scheme [ELS20]. Albeit being aware of the existence of such objects, explicit descriptions are limited to a few instances: smooth quasi-projective schemes over a field [Rou08], Frobenius pushforwards on a compact generator for Noetherian schemes of prime characteristic [${}^{+}$23], singular varieties that admit a resolution of singularities [Lan23, DL24], and objects arising from module categories [DLT23].
Let’s now focus on understanding the Rouquier dimension of $D^{b}_{\operatorname{coh}}(X)$ in the context of a singular variety $X$. In [Lan23, DL24], it was demonstrated that if $\pi\colon\widetilde{X}\to X$ is a resolution of singularities and $G$ is a strong generator for $D^{b}_{\operatorname{coh}}(\widetilde{X})$, then $\mathbb{R}\pi_{\ast}G$ is a strong generator for $D^{b}_{\operatorname{coh}}(X)$. This provides effective methods for explicitly describing strong generators although the associated generation time remains difficult to establish.
Recent attention has been drawn to the case where $X$ is a projective curve [BD11, BDG17b, HH23, BD17] with significant progress made in establishing upper bounds on the Rouquier dimension of its bounded derived category. In particular, [BDG17b] identified an essentially surjective functor $D^{b}_{\operatorname{coh}}(\mathbb{X})\to D^{b}_{\operatorname{coh}}(X)$, where $\mathbb{X}$ is a noncommutative projective curve, yielding a categorical resolution of $X$. In the context of this work, a noncommutative structure sheaf is defined as a pair $(X,\mathcal{A})$, where $X$ is a Noetherian scheme, and $\mathcal{A}$ is a coherent $\mathcal{A}$-algebra [YZ06, BDG17a]. Additional background information can be found in Section 3.1. For alternative approaches to noncommutative algebraic geometry, refer to [Orl16, KR00, Lau03, ${}^{+}$16].
In a closely related context, [ELS20] demonstrated that the bounded derived category of coherent $\mathcal{A}$-modules, denoted $D^{b}_{\operatorname{coh}}(\mathcal{A})$, admits a classical generator whenever $\mathcal{A}$ is a coherent $\mathcal{O}_{X}$-algebra over a Noetherian $J\textrm{-}2$ scheme. Our work is motivated by two key aspects: the preservation of generation along a derived pushforward of a proper surjective morphism and the geometry of a noncommutative structure sheaf. This represents a blend of ideas from [Lan23, DL24, ELS20, BDG17a] and leads us to our first result: an investigation into the preservation of classical generators through the derived pushforward along the canonical morphism from a noncommutative structure sheaf to its underlying scheme. The following is Theorem 3.1.
Theorem A.
Let $X$ be a Noetherian $J\textrm{-}2$ scheme of finite Krull dimension. Suppose $\mathcal{A}$ is a coherent $\mathcal{O}_{X}$-algebra with full support and structure map $\pi\colon\mathcal{O}_{X}\to\mathcal{A}$. If $G$ is a classical generator for $D^{b}_{\operatorname{coh}}(\mathcal{A})$, then $\mathbb{R}\pi_{\ast}G$ is a classical generator for $D^{b}_{\operatorname{coh}}(X)$.
There are interesting examples where Theorem A can be studied, as explored in more detail in Section 4. In particular, these include noncommutative resolutions (NCR), categorical resolution of singularities, and homological projective duality. These examples not only showcase interesting applications of noncommutative structure sheaves but also provide valuable strategies offering upper bounds for Rouquier dimension within the commutative setting. Furthermore, we introduce a construction that yields examples of NCRs, where the obtained bounds indicate that the Rouquier dimension is no larger than the global dimension of a specific endomorphism ring. This is illustrated in Lemma 4.2.3 and discussed further in Remark 4.2.4. The subsequent Corollary 3.2 encapsulates the results presented below.
Corollary B.
Let $R$ be a commutative Noetherian $J\textrm{-}2$ ring of finite Krull dimension and $\pi\colon R\to S$ a finite ring morphism such that $\pi_{\ast}S$ has full support as an $R$-module. If $G$ is a classical generator for $D^{b}_{\operatorname{coh}}(S)$ which is a bounded complex of $(S,S)$-bimodules, then
$\dim D^{b}_{\operatorname{coh}}(R)\leq\operatorname{level}^{\pi_{\ast}G}(R)\cdot\big{(}\dim D^{b}_{\operatorname{coh}}(S)+1\big{)}-1.$
In Corollary B, we may relax the requirement for the classical generator $G$ to be a bounded complex of $(S,S)$-modules, as mentioned in Remark 3.2.10. For those with a geometric inclination, Theorem 3.1 yields an intriguing consequence that is inspired by the exploration of derived categories of Azumaya algebras in [Kuz06], cf. Corollary 3.3.
Corollary C.
Let $X$ be an embeddable variety over a field. Suppose $\mathcal{A}$ is an Azumaya algebra on $X$. If $G$ is a classical generator for $D^{b}_{\operatorname{coh}}(\mathcal{A})$, then
$\dim D^{b}_{\operatorname{coh}}(X)\leq\operatorname{level}^{\pi_{\ast}G}(\mathcal{O}_{X})\cdot(\dim D^{b}_{\operatorname{coh}}(\mathcal{A})+1)-1.$
Initially, we had a similar result for Azumaya algebras over an embeddable variety utilizing homotopy colimits and a form of generation in the unbounded derived category, but the bound was far less sharp. This is discussed in Remark 3.2.12. Our last result provides lower bounds for the Rouquier dimension of the bounded derived category of a coherent $\mathcal{O}_{X}$-algebras on an integral scheme of finite type over a field, cf. Proposition 3.4. It is interesting to note that a similar bound holds for varieties over a field, see [Rou08, Proposition 7.16]. The following result is a special case of a more general statement we prove later, cf. Proposition 3.2.14.
Proposition D.
Let $X$ be an integral scheme of finite type over a field. If $\mathcal{A}$ is a coherent $\mathcal{O}_{X}$-algebra with full support, then the Rouquier dimension of $\dim D^{b}_{\operatorname{coh}}(\mathcal{A})$ is at least $\dim X$.
Acknowledgements.
The authors would like to thank Ryo Takahashi for clarifying Lemma 3.2.8 and Timothy De Deyn for pointing out typos in an earlier draft. Souvik Dey was partially supported by the Charles University Research Center program No. UNCE/24/SCI/022 and a grant GA CR 23-05148S from the Czech Science Foundation. Moreover, the third author would like to personally thank Noah Olander for discussions on homotopy colimits and its relation to big generation, cf. Remark 2.10 for details. In our earlier draft, it was these techniques that were utilized to obtain Corollary 3.3, but recent work of [LO24] have allowed us to obtain sharper bounds for the same statement with a different strategy in the proof.
2. Generation
This section revisits two crucial notions of generation in triangulated categories, with a primary focus on those categories constructed from quasi-coherent sheaves on a Noetherian scheme. For foundational information on triangulated categories and generation, please consult [Kra22, Huy06, BvdB03, Rou08, Nee21, ABIM10, OŠ12a, Let21]. We fix a triangulated category $\mathcal{T}$ with shift functor $[1]\colon\mathcal{T}\to\mathcal{T}$.
Convention 2.1.
Unless otherwise stated, Noetherian ring will mean a right Noetherian and a module will mean a right module.
Notation 2.2.
Let $X$ be a Noetherian scheme.
  • $D_{\operatorname{Qcoh}}(X)$ is the derived category of complexes of $\mathcal{O}_{X}$-modules whose cohomology is quasi-coherent
  • $D^{b}_{\operatorname{coh}}(X)$ is the derived category of bounded complexes of $\mathcal{O}_{X}$-modules whose cohomology is coherent.
By abuse of notation, we let $D^{b}_{\operatorname{coh}}(R)$ denote $D^{b}_{\operatorname{coh}}(\operatorname{Spec}(R))$ if $R$ is a commutative Noetherian ring.
Definition 2.3.
A full triangulated subcategory $\mathcal{S}$ of $\mathcal{T}$ is thick whenever it is closed under retracts. The smallest thick subcategory in $\mathcal{T}$ containing $\mathcal{S}$ is denoted $\langle\mathcal{S}\rangle$. If $\mathcal{S}$ consist of a single object $S$, then we set $\langle\mathcal{S}\rangle=\langle S\rangle$.
Definition 2.4.
Let $\mathcal{S}$ be a subcategory of $\mathcal{T}$.
  1. (1)
    $\langle\mathcal{S}\rangle_{0}$ is the full subcategory consisting of all objects that are isomorphic to the zero object
  2. (2)
    $\langle\mathcal{S}\rangle_{1}$ is the full subcategory containing $G$ closed under shifts and retracts of finite coproducts
  3. (3)
    For $n\geq 2$, $\langle\mathcal{S}\rangle_{n}$ denotes the full subcategory of objects which are retracts of an object $E$ appearing in a distinguished triangle
    $A\to E\to B\to A[1]$
    where $A\in\langle\mathcal{S}\rangle_{n-1}$ and $B\in\langle\mathcal{S}\rangle_{1}$.
Remark 2.5.
From this construction, it may be verified that there is an exhaustive filtration:
$\langle\mathcal{S}\rangle_{0}\subseteq\langle\mathcal{S}\rangle_{1}\subseteq\cdots\subseteq\bigcup_{n=0}^{\infty}\langle\mathcal{S}\rangle_{n}=\langle\mathcal{S}\rangle.$
Definition 2.6.
Let $E,G$ be objects of $\mathcal{T}$. The object $G$ finitely builds $E$ if $E$ belongs to $\langle G\rangle$, and if every object of $\mathcal{T}$ is finitely built by $G$, then we say $G$ is a classical generator. Moreover, if there exists an $n\geq 0$ such that $\langle G\rangle_{n}=\mathcal{T}$, then $G$ is called a strong generator.
Example 2.7.
  1. (1)
    [ELS20, Theorem 4.15] If $X$ is a Noetherian $J\textrm{-}2$, then $D^{b}_{\operatorname{coh}}(X)$ of admits a classical generator.
  2. (2)
    [Aok21, Main Theorem] If $X$ is a Noetherian quasi-excellent separated scheme of finite Krull dimension, then $D^{b}_{\operatorname{coh}}(X)$ admits a strong generator.
  3. (3)
    [${}^{+}$23, Corollary 3.9] If $X$ is a Noetherian $F$-finite scheme, $G$ is a compact generator, and $e\gg 0$, then $F_{\ast}^{e}G$ is a classical generator for $D^{b}_{\operatorname{coh}}(X)$.
Definition 2.8.
Suppose $A,B,C$ are objects of $\mathcal{T}$. If $A$ belongs to $\langle B\rangle$, then we say the level of $A$ with respect to $B$ is the minimal $n$ required such that $A$ belongs to $\langle B\rangle_{n}$. This value is denoted by $\operatorname{level}_{\mathcal{T}}^{B}(A)$. If $C$ is a strong generator for $\mathcal{T}$, then its generation time is the minimal $m$ needed so that the level of an object $A$ in $\mathcal{T}$ with respect to $C$ is at most $m+1$. The smallest integer $d$ such that there exists a strong generator $G$ in $\mathcal{T}$ whose generation time is $d+1$ is called the Rouquier dimension of $\mathcal{T}$, and it is denoted $\dim\mathcal{T}$.
Let us assume that our triangulated category $\mathcal{T}$ admits all small coproducts. It should be noted that the following discussion is utilized only in Corollary 3.3. Fix an object $G$ of $\mathcal{T}$. Let $\overline{\langle G\rangle}$ denote the smallest triangulated subcategory in $\mathcal{T}$ containing $G$ that is closed under small coproducts and retracts. There exists an ascending chain of additive subcategories in $\overline{\langle G\rangle}$:
$\overline{\langle G\rangle}_{0}\subseteq\overline{\langle G\rangle}_{1}\subseteq\cdots\subseteq\overline{\langle G\rangle}$
where each $\overline{\langle G\rangle}_{i}$ is defined as follows
  1. (1)
    $\overline{\langle G\rangle}_{0}$ is the full subcategory consisting of all objects isomorphic to the zero object
  2. (2)
    $\overline{\langle G\rangle}_{1}$ is the smallest full subcategory containing $G$ closed under shifts and retracts of small coproducts
  3. (3)
    For any integer $n\geq 2$, $\overline{\langle G\rangle}_{n}$ is the full subcategory of objects which are retracts of an object $E$ appearing in a distinguished triangle
    $A\to E\to B\to A[1]$
    where $A\in\overline{\langle G\rangle}_{n-1}$ and $B\in\overline{\langle G\rangle}_{1}$.
Remark 2.9.
Let $X$ be a Noetherian scheme. For any $G\in D^{b}_{\operatorname{coh}}(X)$ and integer $n\geq 0$, the following can be deduced from the proofs of Lemma 2.4-2.6 in [Nee21]:
$\overline{\langle G\rangle}_{n}\cap D^{b}_{\operatorname{coh}}(X)=\langle G\rangle_{n}.$
Consequently, if $\overline{\langle G\rangle}_{n}=D_{\operatorname{Qcoh}}(X)$, then $D^{b}_{\operatorname{coh}}(X)=\langle G\rangle_{n}$.
Remark 2.10.
(Olander) Let $X$ be a Noetherian scheme, and choose $E\in D^{b}_{\operatorname{coh}}(X)$. There exists a system $(P_{n},f_{n})$ in $D_{\operatorname{Qcoh}}(X)$, where $P_{n}\in\operatorname{perf}X$ and a distinguished triangle
[…]
where $f$ is given by $1-f_{n}$ in degree $n$, cf. [Sta23, 0DJN]. Hence from Remark 2.9 it follows that if $\operatorname{Perf}(X)\subseteq\langle G\rangle_{n}$, then $D^{b}_{\operatorname{coh}}(X)=\langle G\rangle_{2n}$.
3. Derived descent
In this section, we investigate the preservation of classical generators through the derived pushforward along the canonical morphism from a noncommutative structure sheaf to its underlying scheme. The subsequent is our main result.
Theorem 3.1.
Let $X$ be a Noetherian $J\textrm{-}2$ scheme of finite Krull dimension. Suppose $\mathcal{A}$ is a coherent $\mathcal{O}_{X}$-algebra with full support and structure map $\pi\colon\mathcal{O}_{X}\to\mathcal{A}$. If $G$ is classical generator for $D^{b}_{\operatorname{coh}}(\mathcal{A})$, then $\mathbb{R}\pi_{\ast}G$ is a classical generator for $D^{b}_{\operatorname{coh}}(X)$.
The application of Theorem 3.1 yields immediate consequences, each with distinct algebraic and geometric flavors. These consequences offer strategies coming from noncommutative objects, providing upper bounds for Rouquier dimension within the commutative setting.
Corollary 3.2.
Let $R$ be a commutative Noetherian $J\textrm{-}2$ ring of finite Krull dimension and $\pi\colon R\to S$ a finite ring morphism such that $\pi_{\ast}S$ has full support as an $R$-module. If $G$ is a classical generator for $D^{b}_{\operatorname{coh}}(S)$ which is a bounded complex of $(S,S)$-bimodules, then
$\dim D^{b}_{\operatorname{coh}}(R)\leq\operatorname{level}^{\pi_{\ast}G}(R)\cdot\big{(}\dim D^{b}_{\operatorname{coh}}(S)+1\big{)}-1.$
In Corollary 3.2, our requirement for the classical generator $G$ to be a bounded complex of $(S,S)$-modules is relatively mild and can be further relaxed (see Remark 3.2.10). Next up, we present a purely geometric application inspired by the functorial properties explored for Azumaya varieties and their bounded derived categories in [Kuz06].
Corollary 3.3.
Let $X$ be an embeddable variety over a field. Suppose $\mathcal{A}$ is an Azumaya algebra on $X$. If $G$ is a classical generator for $D^{b}_{\operatorname{coh}}(\mathcal{A})$, then
$\dim D^{b}_{\operatorname{coh}}(X)\leq\operatorname{level}^{\pi_{\ast}G}(\mathcal{O}_{X})\cdot(\dim D^{b}_{\operatorname{coh}}(\mathcal{A})+1)-1.$
For a less sharp bound compared to Corollary 3.3, but whose techniques are interesting in their own right, please refer to Remark 3.2.12 where we utilize homotopy colimits and a form of generation in the unbounded derived category. The last result provides lower bounds for the Rouquier dimension of the bounded derived category of a coherent $\mathcal{O}_{X}$-algebras on an integral scheme of finite type over a field.
Proposition 3.4.
Let $X$ be an integral scheme of finite type over a field. If $\mathcal{A}$ is a coherent $\mathcal{O}_{X}$-algebra with full support, then the Rouquier dimension of $\dim D^{b}_{\operatorname{coh}}(\mathcal{A})$ is at least $\dim X$.
Before delving into the proofs of our results, we take a brief detour to explore the essential background and machinery related to noncommutative structure sheaves.
3.1. Noncommutative structure sheaves
This section draws directly on content found in [YZ06, ELS20, BDG17a]. For further details, the reader is encouraged to refer to these sources; especially [ELS20, Section 4] where background is sourced from. Consider a Noetherian scheme $X$. Let $\mathcal{A}$ be an $\mathcal{O}_{X}$-algebra. The notions of coherent and quasi-coherent $\mathcal{A}$-modules are defined in the obvious way. We denote the respective full subcategories of quasi-coherent and coherent sheaves in the category $\operatorname{Mod}\mathcal{A}$ of $\mathcal{A}$-modules by $\operatorname{Qcoh}\mathcal{A}$ and $\operatorname{coh}\mathcal{A}$. Note that $\operatorname{coh}\mathcal{A}$ is a full abelian subcategory of $\operatorname{Mod}\mathcal{A}$, which is also abelian. Since $\mathcal{A}$ is an $\mathcal{O}_{X}$-algebra, there exists a corresponding structure map $\pi\colon\mathcal{O}_{X}\to\mathcal{A}$. If $E\in\operatorname{Mod}\mathcal{A}$, we can view $E$ as a sheaf of $\mathcal{O}_{X}$-modules by locally restricting scalars via $\pi$. We denote the resulting sheaf of $\mathcal{O}_{X}$-modules as $\pi_{\ast}E$.
Remark 3.1.1.
The following is [ELS20, Lemma 4.1].
  1. (1)
    Assume $\mathcal{A}\in\operatorname{Qcoh}X$. An $\mathcal{A}$-module $E$ is quasi-coherent if, and only if, $\pi_{\ast}E$ is quasi-coherent as an $\mathcal{O}_{X}$-module.
  2. (2)
    Assume $\mathcal{A}\in\operatorname{coh}X$. An $\mathcal{A}$-module $E$ is coherent if, and only if, $\pi_{\ast}E$ is coherent as an $\mathcal{O}_{X}$-module.
Definition 3.1.2.
  1. (1)
    A noncommutative structure sheaf is a pair $(Y,\mathcal{B})$ where $Y$ is a scheme and $\mathcal{B}$ is a quasi-coherent sheaf of $\mathcal{O}_{Y}$-algebras.
  2. (2)
[…]
Example 3.1.3.
If $X$ is a scheme, then the pair $(X,\mathcal{O}_{X})$ is a noncommutative structure sheaf. Assume $\mathcal{A}$ is a quasi-coherent $\mathcal{O}_{X}$-algebra. Consider the morphism of noncommutative structure sheaves
$(1_{X},\pi)\colon(X,\mathcal{A})\to(X,\mathcal{O}_{X}).$
This pair is referred to as the canonical morphism. To simplify notation, we abuse notation to denote it as $\pi$.
Example 3.1.4.
Let $X$ be a variety over a field. Consider a sheaf of Azumaya algebras $\mathcal{A}$ on $X$, i.e., a locally finite free sheaf on $X$ that is étale locally a matrix algebra. This sheaf corresponds to an element in the Brauer group of $X$. For a treatment of the functorial properties of the noncommutative structure sheaf $(X,\mathcal{A})$ and its bounded derived category, refer to [Kuz06, Appendix D]. The pair $(X,\mathcal{A})$ is called a Azumaya variety.
Remark 3.1.5.
Let $\pi\colon Y\to X$ be a morphism of schemes, and set $\mathcal{B}:=\pi^{\ast}\mathcal{A}$
  1. (1)
    Remark 3.1.1 tells us $\mathcal{B}:=\pi^{\ast}\mathcal{A}$ is an $\mathcal{O}_{Y}$-algebra. Moreover, if $\mathcal{A}$ is (quasi-)coherent, then so is $\mathcal{B}$.
  2. (2)
    If $\pi^{\ast}$ and $\pi_{\ast}$ are defined in the usual manner, then we have an adjunction which commutes with the forgetful functor to modules over the underlying schemes:
[…]
  3. (3)
    If $\mathcal{A}$ is quasi-coherent, then the adjunction above restrict to one between quasi-coherent sheaves.
  4. (4)
    If $\mathcal{A}$ is coherent, then $\pi^{\ast}$ restricts to $\pi^{\ast}\colon\operatorname{coh}\mathcal{A}\to\operatorname{coh}\mathcal{B}$.
For further details, please refer to [ELS20, Section 4].
The following Verdier localization sequence will be vital for the proof of Theorem 3.1. Assume $\mathcal{A}$ is a coherent $\mathcal{O}_{X}$-algebra. Denote the derived category of complexes of coherent $\mathcal{A}$-modules by $D_{\operatorname{coh}}(\mathcal{A})$, and $D^{b}_{\operatorname{coh}}(\mathcal{A})$ for the full subcategory of bounded complexes. If $Z$ is a closed subscheme of $X$, we denote by $D^{b}_{\operatorname{coh},Z}(\mathcal{A})$ the full subcategory of $D^{b}_{\operatorname{coh}}(\mathcal{A})$ whose cohomology sheaves are supported in $Z$ when viewed as $\mathcal{O}_{X}$-modules. Note that $D^{b}_{\operatorname{coh},Z}(\mathcal{A})$ is a thick subcategory of $D^{b}_{\operatorname{coh}}(\mathcal{A})$.
Remark 3.1.6.
If $j\colon U\to X$ is an open immersion, then there exists a Verdier localization sequence
$D^{b}_{\operatorname{coh},Z}(\mathcal{A})\to D^{b}_{\operatorname{coh}}(\mathcal{A})\xrightarrow{j^{\ast}}D^{b}_{\operatorname{coh}}(j^{\ast}\mathcal{A})$
where $Z:=X\setminus Y$. This is [ELS20, Theorem 4.4].
Example 3.1.7.
Let $X$ be a Noetherian $J\textrm{-}2$ scheme. If $\mathcal{A}$ is a coherent $\mathcal{O}_{X}$-algebra, then $D^{b}_{\operatorname{coh}}(\mathcal{A})$ admits a classical generator. This is [ELS20, Theorem 4.15].
Remark 3.1.8.
Suppose $i\colon Z\to X$ is a closed immersion. Note $i_{\ast}\colon\operatorname{Mod}Z\to\operatorname{Mod}X$ is exact and preserves coherence. If $\mathcal{A}$ is a coherent $\mathcal{O}_{X}$-module, then $i_{\ast}\colon\operatorname{coh}i^{\ast}\mathcal{A}\to\operatorname{coh}\mathcal{A}$ is well-defined and exact. Hence, this gives an induced (exact) derived functor $i_{\ast}\colon D^{b}_{\operatorname{coh}}(i^{\ast}\mathcal{A})\to D^{b}_{\operatorname{coh}}(\mathcal{A})$. Moreover, if $G$ is a classical generator for $D^{b}_{\operatorname{coh}}(i^{\ast}\mathcal{A})$, then $i_{\ast}G$ is a classical generator for $D^{b}_{\operatorname{coh},Z}(\mathcal{A})$. This is [ELS20, Proposition 4.6].
3.2. Results
To exhibit Theorem 3.1, we start by checking for exactness of the canonical morphism of a noncommutative structure sheaf to its underlying scheme. The following statement might be clear to experts, but we include it for the sake of coherence.
Lemma 3.2.1.
Let $X$ be a Noetherian scheme. If $\mathcal{A}$ is a coherent $\mathcal{O}_{X}$-algebra with structure map $\pi\colon\mathcal{O}_{X}\to\mathcal{A}$, then there exists an exact functor $\pi_{\ast}\colon\operatorname{coh}\mathcal{A}\to\operatorname{coh}X$.
Proof.
If $E\in\operatorname{coh}\mathcal{A}$, then $\pi_{\ast}E\in\operatorname{coh}X$ by Remark 3.1.1. This establishes a well-defined functor $\pi_{\ast}\colon\operatorname{coh}\mathcal{A}\to\operatorname{coh}X$. To verify exactness, consider a short exact sequence in $\operatorname{coh}\mathcal{A}$:
$0\to A\to B\to C\to 0$
This is a local problem on $X$, so we may assume $X=\operatorname{Spec}(R)$ for some commutative Noetherian ring $R$, and $\mathcal{A}$ can be replaced by a finite $R$-algebra $S$. However, $\pi\colon\operatorname{mod}S\to\operatorname{mod}R$ is exact, as the restriction of scalars does not alter the underlying abelian groups. This completes the proof. ∎
Lemma 3.2.2.
Let $X$ be a Noetherian scheme. If $\mathcal{A}$ is a coherent $\mathcal{O}_{X}$-algebra with structure map $\pi\colon\mathcal{O}_{X}\to\mathcal{A}$, then there exists an exact functor $\pi_{\ast}\colon D^{b}_{\operatorname{coh}}(\mathcal{A})\to D^{b}_{\operatorname{coh}}(X)$ of triangulated categories.
Proof.
Lemma 3.2.1 tells us $\pi_{\ast}\colon\operatorname{coh}\mathcal{A}\to\operatorname{coh}X$ is exact on abelian categories, so there exists an induced functor on derived categories $D_{\operatorname{coh}}(\mathcal{A})\to D_{\operatorname{coh}}(X)$. To see $\pi_{\ast}$ preserves bounded cohomology, this follows from the fact that $\pi$ is exact. ∎
Remark 3.2.3.
Let $R$ be a commutative Noetherian ring. Suppose $P,Q\in\operatorname{perf}R$. The following are equivalent: $P$ finitely builds $Q$ if, and only if, $\operatorname{Supp}(Q)\subseteq\operatorname{Supp}(P)$. This is [Nee92, Lemma 1.2] and [Hop87, Theorem 11].
Remark 3.2.4.
Let $X$ be a Noetherian scheme. Suppose that $j\colon U\to X$ is an open subscheme. Choose $E,P\in D^{b}_{\operatorname{coh}}(X)$. If $\mathbb{L}j^{\ast}E\in\langle\mathbb{L}j^{\ast}P\rangle$ in $D^{b}_{\operatorname{coh}}(U)$, then there exists $C\in D^{b}_{\operatorname{coh},Z}(X)$ such that $E\in\langle P\oplus C\rangle$ in $D^{b}_{\operatorname{coh}}(X)$ where $Z=X\setminus U$. This is [Lan23, Lemma 4.1.2].
Remark 3.2.5.
Let $X$ be a Noetherian scheme. If $X=\bigcup^{n}_{i=1}Z_{i}$ denotes the maximal irreducible components and $G_{i}\in D^{b}_{\operatorname{coh}}(Z_{i})$ is a classical generator for each $1\leq i\leq n$, then $\oplus_{i=1}^{n}\mathbb{R}\pi_{i,\ast}G_{i}$ is a classical generator for $D^{b}_{\operatorname{coh}}(X)$ where $\pi_{i}\colon Z_{i}\to X$ is the closed immersion. This can be shown from [DL24, Proposition 3.3].
Proof of Theorem 3.1.
If $X=\emptyset$, there is nothing to check, so without loss of generality, $X$ is non-empty. Note any closed subscheme $Z$ of $X$ is Noetherian $J\textrm{-}2$. By Noetherian induction, we assume the claim holds for any properly contained closed subscheme $Z$ of $X$.
First, consider the case where $X$ is an integral scheme. Then $\operatorname{reg}(X)$ is a nonempty open subscheme of $X$. Let $i\colon U\to X$ be an open immersion, with $U$ an open affine subscheme contained in $\operatorname{reg}(X)$. Fix a classical generator $E$ for $D^{b}_{\operatorname{coh}}(X)$. Then $i^{\ast}\pi_{\ast}G$ is a classical generator for $D^{b}_{\operatorname{coh}}(U)$ by Remark 3.2.3. Thus, $i^{\ast}\pi_{\ast}G$ finitely builds $i^{\ast}E$ in $D^{b}_{\operatorname{coh}}(U)$. Consequently, there exists $A\in D^{b}_{\operatorname{coh}}(Z)$, where $Z$ is a closed subscheme contained in $X\setminus U$, with a closed immersion $j\colon Z\to X$, such that $E$ is finitely built by $\pi_{\ast}G\oplus j_{\ast}A$; cf. Remark 3.2.4. We then demonstrate that $\pi_{\ast}G$ finitely builds $j_{\ast}A$ in $D^{b}_{\operatorname{coh}}(X)$. There exists a commutative diagram of $\mathcal{O}_{X}$-algebras:
[…]
Since $Z$ is a properly contained subscheme of $X$, the induction hypothesis states that if $G^{\prime}$ is a classical generator for $D^{b}_{\operatorname{coh}}(j^{\ast}\mathcal{A})$, then $\theta_{\ast}G^{\prime}$ is a classical generator for $D^{b}_{\operatorname{coh}}(Z)$. Consequently, $(\phi_{2}\circ\theta)_{\ast}G^{\prime}$ finitely builds $j_{\ast}A$ in $D^{b}_{\operatorname{coh}}(X)$. Moreover, since $G$ finitely builds $\phi_{1,\ast}G^{\prime}$ in $D^{b}_{\operatorname{coh}}(\mathcal{A})$, we have that $\pi_{\ast}G$ finitely builds $(\pi\circ\phi_{1})_{\ast}G^{\prime}$ in $D^{b}_{\operatorname{coh}}(X)$. Hence, $\pi_{\ast}G$ finitely builds $j_{\ast}A$ in $D^{b}_{\operatorname{coh}}(X)$ and, consequently, finitely builds $E$.
Lastly, in the general case where $X$ is not an integral domain, a similar argument applies to show that $\pi_{\ast}G$ finitely builds any classical generator for each maximal irreducible component of $X$, cf. Remark 3.2.5. ∎
Remark 3.2.6.
Let $\pi\colon Y\to X$ be a proper surjective morphism of Noetherian schemes. If $G$ is a classical generator for $D^{b}_{\operatorname{coh}}(Y)$, then $\mathbb{R}\pi_{\ast}G$ is a classical generator for $D^{b}_{\operatorname{coh}}(X)$. This is [DL24, Corollary 3.6].
Corollary 3.2.7.
Let $\pi\colon Y\to X$ be a proper surjective morphism of Noetherian $J\textrm{-}2$ schemes of finite Krull dimension. Suppose $\mathcal{A}$ is a coherent $\mathcal{O}_{Y}$-algebra with structure map $\eta\colon\mathcal{O}_{X}\to\mathcal{A}$. If $G$ is a classical generator for $D^{b}_{\operatorname{coh}}(\mathcal{A})$, then $\mathbb{R}(\pi\circ\eta)_{\ast}G$ is a classical generator for $D^{b}_{\operatorname{coh}}(X)$.
Proof.
This is Remark 3.2.6 and Theorem 3.1. ∎
Lemma 3.2.8.
Let $R,S$ be Noetherian rings and $\pi\colon R\to S$ a ring homomorphism. If $E\in D{\operatorname{Qcoh}}(R)$ and $A\in D{\operatorname{Qcoh}}(S)$ is a complex of $(S,S)$-bimodules, then there exists a isomorphism in $D{\operatorname{Qcoh}}(R)$:
$E\overset{\mathbb{L}}{\otimes}_{R}\pi_{\ast}A\cong\pi_{\ast}(\mathbb{L}\pi^{\ast}E\overset{\mathbb{L}}{\otimes}_{S}A).$
Proof.
We can realize $\phi_{\ast}$ and $\mathbb{L}\phi^{\ast}$ as the derived tensor product functors:
$\phi_{\ast}(-)$ $=(-)\overset{\mathbb{L}}{\otimes_{S}}(S_{R}),$
$\mathbb{L}\phi^{\ast}(-)$ $=(-)\overset{\mathbb{L}}{\otimes}_{R}S,$
where $S_{R}$ denotes $S$ as a right $R$-module via $\phi$. There exists an isomorphism in $D{\operatorname{Qcoh}}(R)$:
$\phi_{\ast}(\mathbb{L}\phi^{\ast}E\overset{\mathbb{L}}{\otimes}_{S}A)$ $\cong((E\overset{\mathbb{L}}{\otimes}_{R}S)\overset{\mathbb{L}}{\otimes}_{S}A)\overset{\mathbb{L}}{\otimes}_{S}(S_{R})$
$\cong E\overset{\mathbb{L}}{\otimes}_{R}((S\overset{\mathbb{L}}{\otimes}_{S}A)\overset{\mathbb{L}}{\otimes}_{S}(S_{R}))$
$\cong E\overset{\mathbb{L}}{\otimes}_{R}(A\overset{\mathbb{L}}{\otimes}_{S}(S_{R}))$
$\cong E\overset{\mathbb{L}}{\otimes}_{R}\pi_{\ast}A$
Before obtaining our desired results, we recall a useful statement relating generation time to the number of cones needed to finitely build perfect complexes. The following result is [LO24, Theorem 1.2] which significantly sharpens Remark 2.10.
Theorem 3.2.9 (Lank-Olander).
If $X$ is a Noetherian scheme, $G$ is an object of $D^{b}_{\operatorname{coh}}(X)$, $d$ a non-negative integer, then the following are equivalent:
  1. (1)
    $D^{b}_{\operatorname{coh}}(X)=\langle G\rangle_{d}$
  2. (2)
    $\operatorname{perf}X\subseteq\langle G\rangle_{d}$.
From this result, we now are able to prove Corollary 3.2.
Proof of Corollary 3.2.
If $D^{b}_{\operatorname{coh}}(S)$ has infinite Rouquier dimension, then there is nothing to check, so without loss of generality we may assume this value is finite. Suppose $G$ is a strong generator for $D^{b}_{\operatorname{coh}}(S)$ with generation time is $g$. By Theorem 3.1, $\pi_{\ast}G$ is a classical generator for $D^{b}_{\operatorname{coh}}(R)$. Choose $n\geq 0$ such that $R\in\langle\pi_{\ast}G\rangle_{n}$. According to Lemma 3.2.8, if $P\in\operatorname{perf}R$, the projection formula ensures $P\in\langle\pi_{\ast}(G\overset{\mathbb{L}}{\otimes}\mathbb{L}\pi^{\ast}P)\rangle_{n}$. In particular, this tells us that $\operatorname{perf}R\subseteq\langle\pi_{\ast}G\rangle_{n(g+1)}$. Now, we will utilize this fact in the following. However, the remainder of follows by Theorem 3.2.9. ∎
Remark 3.2.10.
There exists a similar result if, a priori, there exists $E\in D^{b}_{\operatorname{coh}}(S)$ which is a complex of $(S,S)$-bimodules such that $\pi_{\ast}E$ finitely builds $R$ in $D^{b}_{\operatorname{coh}}(R)$. It is worthwhile to note that a similar strategy for showing Corollary 3.2 can be obtained via the proof of [OŠ12b, Theorem 7].
Remark 3.2.11.
Let $X$ be an embeddable variety over a field $k$. Suppose $\mathcal{A}$ is a sheaf of Azumaya algebras on $X$ with structure map $\pi\colon\mathcal{O}_{X}\to\mathcal{A}$. If $E\in D^{-}(\mathcal{A})$ and $G\in D^{-}(X)$, then there exists an isomorphism in $D_{\operatorname{Qcoh}}(X)$:
$\pi_{\ast}(E\overset{\mathbb{L}}{\otimes}\mathbb{L}\pi^{\ast}G)\cong\pi_{\ast}E\overset{\mathbb{L}}{\otimes}G.$
This is [Kuz06, Lemma 2.12].
Proof of Corollary 3.3.
If $D^{b}_{\operatorname{coh}}(\mathcal{A})$ has infinite Rouquier dimension, then there is nothing to show; assume this value is finite. Theorem 3.1 ensures that $\pi_{\ast}G$ is a strong generator for $D^{b}_{\operatorname{coh}}(X)$. Let $P\in\operatorname{perf}X$. There exists $n\geq 0$ such that $\mathcal{O}_{X}\in\langle\pi_{\ast}G\rangle_{n}$. By Remark 3.2.11, we observe that
$P$ $\in\langle\pi_{\ast}G\overset{\mathbb{L}}{\otimes}P\rangle_{n}$
$=\langle\pi_{\ast}(G\overset{\mathbb{L}}{\otimes}\mathbb{L}\pi^{\ast}P)\rangle_{n}.$
This implies $\operatorname{perf}X\subseteq\langle\pi_{\ast}G\rangle_{(g+1)n}$, where $g$ is the generation time. The desired claim follows by Theorem 3.2.9. ∎
Remark 3.2.12.
Let us keep the notation of Corollary 3.3. The following observation provides a weaker upper bound on Rouquier dimesion, but its techniques are worthwhile to write out. If $D^{b}_{\operatorname{coh}}(\mathcal{A})$ has infinite Rouquier dimension, then there is nothing to show; assume this value is finite. Theorem 3.1 ensures that $\pi_{\ast}G$ is a strong generator for $D^{b}_{\operatorname{coh}}(X)$. Let $P\in\operatorname{perf}X$. There exists $n\geq 0$ such that $\mathcal{O}_{X}\in\langle\pi_{\ast}G\rangle_{n}$. By Remark 3.2.11, we observe that
$P$ $\in\langle\pi_{\ast}G\overset{\mathbb{L}}{\otimes}P\rangle_{n}$
$=\langle\pi_{\ast}(G\overset{\mathbb{L}}{\otimes}\mathbb{L}\pi^{\ast}P)\rangle_{n}.$
This implies $\operatorname{perf}X\subseteq\langle\pi_{\ast}G\rangle_{(g+1)n}$, where $g$ is the generation time. Hence, Remark 2.9 ensures $E\in\langle\pi_{\ast}G\rangle_{2(g+1)n}$, and the claim follows.
Proposition 3.2.13.
Let $X$ be a Noetherian $J\textrm{-}2$ scheme, and $\mathcal{A}$ be a coherent $\mathcal{O}_{X}$-algebra. If there exists an open affine immersion $i\colon U\to X$ such that $\mathcal{O}_{U}\in\langle i^{\ast}\mathcal{A}\rangle_{1}$ and $U$ is nonempty, then the Rouquier dimension of $D^{b}_{\operatorname{coh}}(\mathcal{A})$ is at least that of $D^{b}_{\operatorname{coh}}(U)$.
Proof.
Let $i\colon U\to X$ be an open immersion where $U=\operatorname{Spec}(R)$ is a nonempty affine scheme. Consider the Verdier localization $i^{\ast}\colon D^{b}_{\operatorname{coh}}(\mathcal{A})\to D^{b}_{\operatorname{coh}}(i^{\ast}\mathcal{A})$. This tells us that the Rouquier dimension of $D^{b}_{\operatorname{coh}}(\mathcal{A})$ is at least that of $D^{b}_{\operatorname{coh}}(i^{\ast}\mathcal{A})$. Since $U$ is affine and $\mathcal{O}_{U}\in\langle i^{\ast}\mathcal{A}\rangle_{1}$, Corollary 3.2 tells us that the Rouquier dimension of $D^{b}_{\operatorname{coh}}(i^{\ast}\mathcal{A})$ is at least that of $D^{b}_{\operatorname{coh}}(U)$, and this completes the proof. ∎
Proposition 3.2.14.
Let $X$ be an irreducible, Jacobson, catenary, $J\textrm{-}2$, and Noetherian scheme of finite Krull dimension. If $\mathcal{A}$ is a coherent $\mathcal{A}$-algebra, then the Rouquier dimension of $D^{b}_{\operatorname{coh}}(X)$ is at least the Krull dimension of $X$.
Proof.
The hypothesus on $X$ tells us the Krull dimension of any nonempty open affine subscheme of $X$ coincides with that of $X$, cf. [Sta23, Tag 0DRT]. The regular locus of $X$ is nonempty and open, so we can find an open affine immersion $i\colon U\to X$ where $U$ is nonempty. The locus of points $p$ where $\mathcal{A}_{p}$ finitely builds $\mathcal{O}_{X,p}$ in at most one cone in $D^{b}_{\operatorname{coh}}(\mathcal{O}_{X,p})$ is open in $U$, cf. [Let21, Proposition 3.5]. Choose a basic affine open $V=\operatorname{Spec}(R)$ in this locus of points, and note that $\mathcal{A}_{p}$ finitely builds $\mathcal{O}_{X,p}$ in at most one cone in $D^{b}_{\operatorname{coh}}(\mathcal{O}_{X,p})$ for each $p\in V$. Let $j\colon V\to X$ denote the open affine immersion. Then $j^{\ast}\mathcal{A}$ finitely builds $\mathcal{O}_{V}$ in at most one cone in $D^{b}_{\operatorname{coh}}(V)$, cf. [Let21, Corollary 3.4]. Therefore, the desired claim follows immediately by Proposition 3.2.13. ∎
4. Examples
This section presents a compilation of intriguing examples exhibiting instances where Theorem 3.1 can be observed in both algebraic and geometric contexts. It is important to note that this compilation is not exhaustive and represents only a partial exploration of possible applications.
4.1. Algebras of finite type
Let $k$ be a field, and suppose $S$ is a Noetherian ring whose center $R:=Z(S)$ is of finite type over $k$. Assume that $S$ is finite over $R$. Example 3.1.7 tells us that $D^{b}_{\operatorname{coh}}(S)$ admits a classical generator. Note that $D^{b}_{\operatorname{coh}}(S)$ is considered smooth over $k$ in the sense of dg-categories, as discussed in Theorem 5.1 and Definition 2.3 of [ELS20]. From Lemma 3.5 and Lemma 3.6 in [Lun10], it follows that $D^{b}_{\operatorname{coh}}(S)$ admits a strong generator. Further details can be found in [ELS20, Remark 2.6].
4.2. Noncommutative resolutions
The forthcoming examples pertain to noncommutative commutative resolutions. This concept has a rich history, and although not exhaustively listed, key references include [VdB04, SVdB08, ŠVdB17, FMS19, DITV15, IW13]. For our discussion below, we adopt a moderately broad notion introduced in [DFI15]. Consider a commutative Noetherian ring $R$. A noncommutative resolution (NCR), as defined in [DFI15, Definition 2.5], is a finitely generated $R$-module $M$ such that $\operatorname{Supp}(M)=\operatorname{Spec}(R)$ and $\operatorname{End}_{R}(M)$ has finite global dimension. Moreover, if $R\in\operatorname{add}(M)$, then $M$ is termed a generator NCR. It is important to compare the differences in definition for an NCR in the literature. For instance, [DITV15, Definition 1.1] requires the $R$-module to be faithful, whereas [DFI15, Definition 2.5] only requires $M$ to have full support. Moreover, it is of interest to compare these examples to statement such as [IT16, Corollary 6.8] regarding properties on quotients by minimal prime ideals.
Remark 4.2.1.
Suppose $R$ is a commutative Noetherian ring. If $R$ has a generator NCR given by a finitely generated $R$-module $M$, then the Rouquier dimension of $D^{b}_{\operatorname{coh}}(R)$ is at most the global dimension of $\operatorname{End}_{R}(M)$. This follows from Corollary 3.2.
Example 4.2.2.
Any complete local reduced Noetherian ring $R$ of Krull dimension one admits a generator an NCR [DFI15, Example 2.7]. Notably, the existence of a generator NCR implies that the endomorphism ring of $R$ must have global dimension one, which, in turn, implies $R$ is a regular ring [DFI15, Proposition 3.2]. Furthermore, when $R$ is complete, it is possible to construct the generator NCR in such a way that the global dimension of its corresponding endomorphism ring is at most the multiplicity of $R$ [Leu07, Proposition 1.5].
The following statement illustrates a construction for a module $M\in\operatorname{mod}\Lambda$ satisfying the property that $\Lambda$ belongs to $\operatorname{add}_{\Lambda}(M)$ and the endomorphism ring $\operatorname{End}_{\Lambda}(M)$ has a finite global dimension. Recall a Noether algebra is a Noetherian ring $\Lambda$ that is finite over its center $\Lambda^{c}$. We denote a syzygy of module $M$ in $\operatorname{mod}\Lambda$ by $\Omega_{\Lambda}^{1}(M)$, and note that higher syzygies can be defined inductively.
Lemma 4.2.3.
Consider a Noether algebra $\Lambda$. Let $\mathcal{X}$ be a subcategory of $\operatorname{mod}\Lambda$ that is closed under finite direct sums. For any integer $n\geq 0$ and $r\geq 0$, if $M$ is an object in $\mathcal{X}$, then the following assertions hold:
  1. (1)
    Assume that $\mathcal{X}$ is closed under $r$-kernels as defined in [MT22, Definition 3.1]. For any $X\in\mathcal{X}$, if there exists an exact sequence
    $0\to M_{n}\to\cdots\to M_{0}\to X\to 0$
    with $M_{i}\in\operatorname{add}_{\Lambda}(M)$ such that the induced sequence
    $0\to\operatorname{Hom}_{\Lambda}(M,M_{n})\to\cdots\to\operatorname{Hom}_{\Lambda}(M,M_{0})\to\operatorname{Hom}_{\Lambda}(M,X)\to 0$
    is also exact, then the global dimension of $\operatorname{End}_{\Lambda}(M)$ is at most $n+r+3$.
  2. (2)
    Assume that $\mathcal{X}$ is closed under extensions. For any $N\in\Omega_{\Lambda}^{1}(\mathcal{X})$ and $L\in\mathcal{X}$, if there exists a short exact sequence
    $0\to M\to N\to L\to 0,$
    then $M\in\Omega_{\Lambda}^{1}(\mathcal{X})$. Moreover, if $\Omega_{\Lambda}^{1}(\mathcal{X})\subseteq\mathcal{X}$, then $\Omega_{\Lambda}^{1}(\mathcal{X})$ is closed under direct summands.
  3. (3)
    Assume that $\mathcal{X}$ is closed under extensions, $r$-kernels, and $\Lambda\in\mathcal{X}$. If $\Omega_{\Lambda}^{1}(\mathcal{X})\subseteq\operatorname{add}_{\Lambda}(M)$ for some $M\in\Omega_{\Lambda}^{1}(\mathcal{X})$, then $M$ is a generator NCR, and the global dimension of $\operatorname{End}_{\Lambda}(M)$ is at most $r+4$.
Proof.
[…]
$0\to X\to N_{r-1}\to\cdots\to N_{0}\to L_{1}\to L_{0},$
where $X\in\operatorname{mod}\Lambda$. By [MT22, Remark 3.2(4)], $\operatorname{add}_{\Lambda}(M)\subseteq\mathcal{X}$, implying $N_{i}\in\mathcal{X}$ for each $1\leq i\leq r$ and $L_{1},L_{0}\in\mathcal{X}$. As $\mathcal{X}$ is closed under $r$-kernels, we deduce that $X\in\mathcal{X}$. Given our hypothesis, there exists an exact sequence
$0\to M_{n}\to\cdots\to M_{0}\to X\to 0$
with $M_{i}\in\operatorname{add}_{\Lambda}(M)$, ensuring that
$0\to\operatorname{Hom}_{\Lambda}(M,M_{n})\to\cdots\to\operatorname{Hom}_{\Lambda}(M,M_{0})\to\operatorname{Hom}_{\Lambda}(M,X)\to 0$
is also exact. Splicing with the exact sequence
$0\to\operatorname{Hom}_{\Lambda}(M,X)\to\operatorname{Hom}_{\Lambda}(M,N_{r-1})\to\cdots\to\operatorname{Hom}_{\Lambda}(M,N_{0})\to\operatorname{Hom}_{\Lambda}(M,N)\to 0,$
we obtain the following
$0\to$ $\operatorname{Hom}_{\Lambda}(M,M_{n})\to\cdots\to\operatorname{Hom}_{\Lambda}(M,M_{0})$
$\to\operatorname{Hom}_{\Lambda}(M,N_{r-1})\to\cdots\to\operatorname{Hom}_{\Lambda}(M,N_{0})\to\operatorname{Hom}_{\Lambda}(M,N)\to 0$
where each $M_{i},N_{j}\in\operatorname{add}_{\Lambda}(M)$, hence each $\operatorname{Hom}_{\Lambda}(M,M_{i}),\operatorname{Hom}_{\Lambda}(M,_{j})\in\operatorname{proj}T$. Since applying $\operatorname{Hom}_{\Lambda}(M,-)$ to the exact sequence $0\to N\to L_{1}\to L_{0}$ also gives us an exact sequence
$0\to\operatorname{Hom}_{\Lambda}(M,N)\to P_{1}\to P_{0}\to Y\to 0,$
we conclude $\operatorname{pdim}_{T}Y\leq n+r+3$. Since $Y\in\operatorname{mod}T$ is arbitrary, this finishes the claim.
(2) The validity of the latter part of the assertion is evident from the former, achieved by expressing the split short exact sequence corresponding into a direct sum. To establish the first part of the claim, under the given hypothesis, consider the exact sequence $0\to N\to P\to N^{\prime}\to 0$ for some $P\in\operatorname{proj}\Lambda$ and $N^{\prime}\in\mathcal{X}$. The desired conclusion can then be derived from the following pushout diagram:
[…]
remembering that $U\in\mathcal{X}$ because the subcategory $\mathcal{X}$ is extension-closed.
(3) We establish $\Omega_{\Lambda}^{1}(\mathcal{X})\subseteq(\mathcal{X})$ as per [MT22, Remark 3.2(4)]. Consequently, by claim (2), $\Omega_{\Lambda}^{1}(\mathcal{X})$ is closed under direct summands, leading to $\Omega_{\Lambda}^{1}(\mathcal{X})=\operatorname{add}_{\Lambda}(M)$. In particular, this implies that $M$ is a generator NCR. Let $T=\operatorname{End}_{\Lambda}(M)$, and choose $X\in\mathcal{X}$. We know that $\operatorname{Hom}_{\Lambda}(M,X)\in\operatorname{mod}T$. Consider a map $f\colon M_{0}\to X$, where $M_{0}\in\operatorname{add}_{\Lambda}(M)$, and observe that the sequence
$\operatorname{Hom}_{\Lambda}(M,M_{0})\xrightarrow{\operatorname{Hom}_{\Lambda}(M,f)}\operatorname{Hom}_{\Lambda}(M,X)\to 0$
forms a $T$-projective cover of $\operatorname{Hom}_{\Lambda}(M,X)$. Since $M$ is a generator NCR, it follows that $f$ is surjective. As $M_{0}\in\Omega_{\Lambda}^{1}(\mathcal{X})$, we deduce that $\ker f\in\Omega_{\Lambda}^{1}(\mathcal{X})$ by claim (2). Consequently, the sequence
$\sigma\colon 0\to\ker f\to M_{0}\to X\to 0$
is exact, where $\ker f,M_{0}\in\operatorname{add}_{\Lambda}(M)$, and $\operatorname{Hom}_{\Lambda}(M,\sigma)$ is also exact. As $X\in\mathcal{X}$ was chosen arbitrarily, we have demonstrated that the hypotheses of property (1) are satisfied with $n=1$. Therefore, the global dimension of $\operatorname{End}_{\Lambda}(M)$ is at most $r+4$. ∎
Remark 4.2.4.
To apply the above construction in a down-to-earth scenario, consider a Cohen-Macaulay ring $\Lambda$ with Krull dimension $r+1$. Note $\Lambda$ is commutative. In this case, let $\mathcal{X}:=\operatorname{CM}(\Lambda)$, which includes $\Lambda$ and is closed under extensions and $r$-kernels. Notably, if $R$ is a normal domain with Krull dimension two and $\operatorname{Spec}(R)$ has rational singularities, then, as per [IW10, Theorem 3.6(a)] and the above construction, $R$ possesses a generator NCR. Conversely, for a local, excellent, Henselian domain $R$ with an algebraically closed residue field and admitting a generator NCR, the conclusion holds that $\operatorname{Spec}(R)$ has rational singularities, cf. [DITV15, Corollary 3.3].
Example 4.2.5.
Let $R$ be a local Cohen-Macaulay ring of finite Cohen-Macaulay representation type. There exists a generator NCR $M\in\operatorname{mod}R$ such that the global dimension of $\operatorname{End}_{R}(M)$ is at most $\max\{2,\dim R\}$. This is [Leu07, Theorem 2.1].
4.3. Categorical resolutions
There is a homological construction, called categorical resolutions, that emulates the functorial behavior observed in the derived category of a variety with rational singularities. A substantial body of work has been devoted to this pursuit, and the following list of key references is not complete: [KL15, Lun10, Kuz07, BBF16, DD23, Lun12, KS22, Abu16]. Let $X$ be a separated Noetherian scheme of finite type over a field. A categorical resolution of $X$ is a smooth cocomplete compactly generated triangulated category $\mathcal{T}$ with a pair of adjoint functors $\pi_{\ast}\colon\mathcal{T}\to D_{\operatorname{Qcoh}}(X)$ and $\pi^{\ast}\colon D_{\operatorname{Qcoh}}(X)\to\mathcal{T}$ such that
  1. (1)
    $\pi_{\ast}\pi^{\ast}\cong 1_{D_{\operatorname{Qcoh}}(X)}$
  2. (2)
    $\pi^{\ast}$ and $\pi_{\ast}$ commute with small coproducts
  3. (3)
    $\pi^{\ast}(\mathcal{T}^{c})$ is contained in $D^{b}_{\operatorname{coh}}(X)$ where $\mathcal{T}^{c}$ is the subcategory of compact objects in $\mathcal{T}$.
There is a variant for a categorical resolution of $D^{b}_{\operatorname{coh}}(X)$, cf. [Kuz07, Definition 3.2], but this is not treated here.
Example 4.3.1.
Let $X$ be a reduced separated Noetherian one-dimensional scheme that is of finite type and projective over a field $k$. Suppose $\nu\colon\widetilde{X}\to X$ is the normalization of $X$. There exists an inductively constructed collection of $\mathcal{O}_{X}$-modules:
$\mathcal{O}_{1}:=\mathcal{O}_{X}\subseteq\mathcal{O}_{2}\subseteq\mathcal{O}_{3}\subseteq\cdots\subseteq\mathcal{O}_{n}:=\nu_{\ast}\mathcal{O}_{\widetilde{X}}$
[…]
Example 4.3.2.
Consider a separated Noetherian scheme of finite type over a field $k$. Suppose there exists an non-negative integer $n$ and ideal sheaf $\mathcal{I}$ on $X$ such that $\mathcal{I}^{n}=(0)$. Consider the coherent sheaf of $\mathcal{O}_{X}$-algebras
[…]
Assume the closed subscheme $Z$ of $S$ corresponding to the ideal $\mathcal{I}$ is smooth. There exists a semiorthogonal decomposition
$D^{b}_{\operatorname{coh}}(\mathcal{A})=\langle D^{b}_{\operatorname{coh}}(Z),\cdots,D^{b}_{\operatorname{coh}}(Z)\rangle$
where the number of components is equal to $n$. This is [KL15, Corollary 5.15]. Moreover, $D^{b}_{\operatorname{coh}}(\mathcal{A})$ is a categorical resolution of $D^{b}_{\operatorname{coh}}(Z)$, cf. [KL15, Theorem 5.23].
4.4. Homological projective duality
There is a triangulated framework, due to [Kuz06], that enables a rich comparison of the bounded derived categories of two projectively dual varieties which is called homological projective duality (HPD). Informally, HPD states the derived categories of linear sections of a smooth projective mapping to a projective space $X\to\mathbb{P}(V)$ are governed by a (noncommutative) variety $Y\to\mathbb{P}(V^{\vee})$ over the dual projective space, where $Y$ is called the HPD dual of $X$. Recently, HPD has been generalized to broader (noncommutative) contexts [Per19, KP21]. See [Kuz14, Tho18] for survey and additional background. Let $k$ be a field.
Example 4.4.1.
Let $U,V$ be vector spaces over $k$ respectively with dimensions $m,n$. Assume $n\geq m$. Set $W:=U\otimes_{k}V$. Choose an integer $11\leq r\leq m-$ 1. Define $\mathcal{L}^{r}:=\mathcal{L}^{r}_{m,n}$ to be the variety of matrices $M\colon V\to U^{\vee}$ in $\mathbb{P}(W)$ cut by the minors of size $r+1$ of the matrix of indeterminates $(x_{ij,})_{i,j=1}^{m,n}$. Consider the Grassmannian variety $\mathbb{G}:=\mathbb{G}(U,r)$ of $r$-dimensional quotient spaces of $U$. Denote the quotient bundle and tautological subbundle over $\mathbb{G}$ of respective rank $r$ and $m-r$ by $\mathcal{D}$ and $\mathcal{U}$. The Euler sequences gives us
$0\to\mathcal{U}\to U\otimes_{k}\mathcal{O}_{\mathbb{G}}\to\mathcal{D}\to 0$
[…]
Example 4.4.2.
Suppose $V$ is a $k$-vector space of odd dimension $v$. For each even number $0\leq 2q<v$, we have a Pfaffian variety $\operatorname{Pf}_{q}$ contained in $\mathbb{P}(\wedge^{2}V^{\vee})$ consisting of all $2$-forms on $V$ whose rank is at most $2q$. This variety is not a complete intersection and usually is singular. The projective dual of $\operatorname{Pf}_{q}$ is another Pfaffian variety whose locus is the subspace of $\mathbb{P}(\wedge^{2}V)$ consisting of bivectors of rank at most $2s$ where $2s=v-1-2q$. There are noncommutative crepant resolutions for $\operatorname{Pf}_{s}$ and $\operatorname{Pf}_{q}$ respectively denoted by $A$ and $B$, see [ŠVdB17, Section 6]. Note this gives noncommutative structure sheaves $(\operatorname{Pf}_{s},A)$ and $(\operatorname{Pf}_{q},B)$. These are homological projective duals of one another, cf. [RS19]. Let $L$ be a generic subspace of $\wedge^{2}V^{\vee}$ with dimension $sv$. Moreover, the sheaves $A,B$ restrict to give noncommutative crepant resolutions of the varieties $\operatorname{Pf}_{s}\cap\mathbb{P}(L^{\perp}),\operatorname{Pf}_{q}\cap\mathbb{P}(L)$ such that $D^{b}_{\operatorname{coh}}(A|_{\mathbb{P}(L^{\perp})})$ and $D^{b}_{\operatorname{coh}}(B|_{\mathbb{P}(L)})$ are equivalent categories. This is [RS19, Theorem 1.1]. The reader is encouraged to see [RS19] for further details; including the case of even-dimensional vector spaces.